An airplane has a total of 152 seats. The number of coach-class seats is five more than six times the number of first-class seats. How many of each type of seat are there on the plane? Write two equations using c to represent coach-class seats and f to represent first-class seats. Solve using elimination or substitution
Answer by mananth(16949)

c+f=152.............(1)
c= 5+6f
c-6f=5..............(2)
c + 1 f = 152 .............1
c -6 f = 5 .............2
Eliminate y
multiply (1)by 6
Multiply (2) by 1
6 c 6 f = 912
1 c -6 f = 5
Add the two equations
7 c = 917
/ 7
c = 131
plug value of x in (1)
1 c + f = 152
131 + f = 152
f = 152-131
f = 21
f = 21